# Xiaofeng Yang, Yuhui Zhang, Yifeng Zheng, Xiuqi Huang, Minfeng Zhuparallel-$\parallel$, Yuxin Ma, Bin Cui§, Wei Chen

### v2

@@ Abstract
Business intelligence (BI) transforms large volumes of data within modern organizations into actionable insights for informed decision-making. Recently, large language model (LLM)-based agents have streamlined the BI workflow by automatically performing task planning, reasoning, and actions in executable environments based on natural language (NL) queries. However, existing approaches primarily focus on individual BI tasks such as NL2SQL and NL2VIS. The fragmentation of tasks across different data roles and tools lead to inefficiencies and potential errors due to the iterative and collaborative nature of BI. In this paper, we introduce DataLab, a unified BI platform that integrates a one-stop LLM-based agent framework with an augmented computational notebook interface. DataLab supports a wide range of BI tasks for different data roles by seamlessly combining LLM assistance with user customization within a single environment. To achieve this unification, we design a domain knowledge incorporation module tailored for enterprise-specific BI tasks, an inter-agent communication mechanism to facilitate information sharing across the BI workflow, and a cell-based context management strategy to enhance context utilization efficiency in BI notebooks. Extensive experiments demonstrate that DataLab achieves state-of-the-art performance on various BI tasks across popular research benchmarks. Moreover, DataLab maintains high effectiveness and efficiency on real-world datasets from Tencent, achieving up to a 58.58% increase in accuracy and a 61.65% reduction in token cost on enterprise-specific BI tasks.
@@ I Introduction
Recent advancements in autonomous agents powered by large language models (LLMs) [3] offer the potential to streamline the BI workflow (Figure 1). By receiving instructions in natural language (NL), LLM-based agents can perform task planning, reasoning, and actions in executable environments. This can significantly reduce the complexity of many BI tasks, such as code generation [4], text-to-visualization translation [5], and automated insight discovery [6].
@@
Figure 1: General workflow of LLM-based agents for BI tasks.
C1: Lack of domain knowledge incorporation. Existing studies usually leverage clean and synthesized research benchmarks to build and evaluate agents [10]. However, BI tasks typically involve large and dirty real-world datasets with many ambiguities [11]. For example, column names in business data tables may have unclear semantic meanings [12], and user queries often include enterprise-specific jargon [10]. To mitigate these issues, incorporating extensive domain knowledge is essential to enhance agents’ understanding of input data and improve their performance on practical BI tasks. While some approaches adopt fine-tuning [13] or continued pre-training [14] to augment agents’ domain-specific capabilities, acquiring the necessary large and up-to-date training data corpora remains challenging in BI scenarios due to their complexity and dynamic nature.
C2: Insufficient information sharing across tasks. Different tasks are typically managed by corresponding LLM-based agents to achieve optimal performance [15]. As a complex BI query may encompass multiple tasks, information sharing among the involved agents is critical. For example, the data retrieved by a SQL writing agent must be accurately conveyed to a chart generation agent. Therefore, an effective and efficient inter-agent communication mechanism is essential to align their understanding of the overall analysis goals, current data context, and executed actions. However, many existing multi-agent frameworks, such as ChatDev [16] and CAMEL [17], use unstructured natural language for communication. This can lead to distortions [18] and is inadequate for handling the complexity of BI tasks, which commonly involve diverse information types (e.g., data, charts, texts).
C3: Demand for adaptive LLM context management. LLM-based agents depend on their context windows (i.e., limited input tokens for NL understanding, reasoning, and generation) to complete tasks. Necessary contexts must be provided to ensure a successful and seamless workflow. Meanwhile, in a unified BI platform, vast amounts of multi-modal contexts (e.g., code snippets and their execution results, charts and their specifications) are intertwined and often relate to diverse data tables. Obviously, only relevant portions of these contexts are pertinent to specific tasks and should be selectively provided to the agents [19]. Therefore, adaptive context management tailored to prior states and current user needs is crucial for maintaining system efficiency and cost-effectiveness.
@@
To improve agents’ performance on enterprise-specific BI tasks (for C1), we develop a Domain Knowledge Incorporation module, a systematic approach for automated knowledge generation, organization, and utilization. It leverages data processing scripts (e.g., Python code, SQL queries) within the entire enterprise to extract the knowledge associated with databases/tables/columns, thereby uncovering their common usage patterns. To facilitate information sharing across different tasks (for C2), we design an Inter-Agent Communication module, a structured mechanism that goes beyond pure NL to enhance the information representation capabilities of agents. It also formulates the information sharing process among agents with a finite state machine (FSM) for a more controlled and efficient flow of communication. Finally, to manage LLM contexts within multi-modal notebooks (for C3), we propose a Cell-based Context Management module that represents inter-cell dependencies using directed acyclic graphs (DAGs). These dependency graphs are dynamically updated in response to user modifications. This enables the adaptive selection of pertinent contexts based on specific task requirements, thereby enhancing agents’ context utilization efficiency.
@@
  • We extensively evaluate DataLab on both research benchmarks and real-world business datasets from Tencent, demonstrating its performance on various BI tasks.
@@ II Background
In this section, we briefly introduce the core stages and roles in a modern BI workflow. We then provide an overview of LLM-based agents, focusing on their applications in BI tasks.
@@ II-A BI Workflow
The BI workflow encompasses several critical stages, namely data collection, storage, preparation, analysis, and visualization. Data Collection [20] refers to the initial step of gathering raw data from various sources like databases and spreadsheets. Once gathered, Data Storage [21] typically uses data warehouses or data lakes to organize the collected data for efficient retrieval. This may also involve combining data into a unified format (i.e., data integration). Subsequently, Data Preparation [22] ensures data consistency, correctness, and quality. This usually includes cleaning, structuring, and enriching raw data into a format suitable for further analysis. Following preparation, Data Analysis [23] applies statistical and analytical techniques to extract insights, aiming to uncover patterns, trends, and correlations in the data. Finally, Data Visualization [5] presents analyzed data in visual formats like charts, graphs, and dashboards, which makes complex data easier to understand and interpret for decision-makers.
The data roles involved in the BI workflow are specific to different organizations. Among them, data engineers, scientists, and analysts are usually indispensable. Data Engineers are tasked with data collection, storage, and preparation, constructing and administering data pipelines to ensure that data is accurately cleansed and structured for subsequent analysis. They typically use SQL and Python for data processing, and use cloud computing platforms like AWS for data storage and ingestion. Data Scientists engage in data preparation and analysis, applying advanced statistical and machine learning methodologies to extract insights and forecast trends from intricate datasets. They are familiar with Python/R and popular data science libraries like Pandas. Data Analysts concentrate on data analysis and visualization, analyzing data to discern patterns and conveying findings through detailed visual reports and dashboards. They use SQL to query data, and rely on BI platforms like Tableau to perform and share their analyses.
@@ II-B LLM-based Agents for BI
LLM-based agents are autonomous systems powered by LLMs that can perceive environments, execute tasks, make decisions, and interact with users in complex contexts [3]. These agents comprise profiling, memory, planning, and action modules, which respectively define the agent’s role, facilitate operations in dynamic environments through recall and future action planning, and convert decisions into outputs [24]. In BI scenarios, agents receive users’ NL queries and then perform data preparation, analysis, and visualization. By interpreting execution results, they can iteratively complete many BI tasks such as data transformation [25] and insight generation [6]. Moreover, equipped with vision language models (VLMs) [26], agents can even generate GUI operations for enterprise applications (e.g., BigQuery, Airflow) [2], which further augments their capabilities for more complex BI workflows. Below, we list some typical tasks that can be streamlined by LLM-based agents at each BI stage:
  • Data Collection: Table Generation [27], Table Augmentation [28], Table Summarization [29].
  • Data Storage: Data Warehousing [2], Data Integration [30], Data Orchestration [2].
  • Data Preparation: NL2SQL [31], NL2DSCode [4].
  • Data Analysis: NL2Insight [32], Table Q&A [33].
  • Data Visualization: NL2VIS [5], Chart Q&A [34].
@@ III Overview
  • LLM-based Agent Framework. In DataLab, multiple agents are designed for different BI tasks based on user requirements. To achieve this, we first identify several common BI procedures and abstract them into data tools that can be called upon by agents during inference. Example tools include a Python sandbox for code execution and a Vega-Lite environment for visualization rendering. Accompanied by other auxiliary components like memory modules, each BI agent is represented as a DAG for high flexibility and easy extensibility. Within the DAG, nodes depict reusable components (e.g., LLM APIs, tools) and edges depict their connections (e.g., file transfer across tools). Figure 3 illustrates an example agent workflow for NL2VIS. Additionally, we add a proxy agent to the framework, which serves as a hub to directly interact with users and allocate tasks to each specialized agent based on user queries. These agents collaborate with each other to complete a wide array of tasks for various data roles throughout the BI workflow.
  • Computational Notebook Interface. DataLab’s notebook interface (Figure 4) serves as a unified, interactive, and collaborative environment for different data roles to complete their specialized tasks. To achieve this, we augment JupyterLab (a widely used notebook interface) to support (1) multi-language cells and (2) on-the-fly LLM assistance. First, DataLab wrangles SQL, Python/PySpark, Markdown, and Chart cells altogether, allowing both technical and non-technical users to easily adopt their familiar workflows on the notebook. Going beyond traditional notebooks that only support Python and Markdown, DataLab notebooks directly connect to backend databases for SQL query execution, and feature GUI-based dashboards [35] similar to Tableau for visualization authoring. Second, we integrate our LLM-based agent framework seamlessly into each notebook cell. Users can get LLM assistance both at notebook- and cell-level. Specifically, users toggle an input box and type their analytic queries, which are then processed by the agents in our framework. These agents can create new cells or modify existing ones in the notebook. Users can subsequently examine the results and make further customizations flexibly.
DataLab Workflow. Upon receiving an NL query and the associated dataset, DataLab initially analyzes the dataset and interprets the query, incorporating domain knowledge [Uncaptioned image] before feeding them into LLMs. Then, DataLab leverages various agents to complete the task, which may involve information sharing with each other through a structured communication mechanism [Uncaptioned image]. Subsequently, the corresponding result will be presented in the notebook. Users can either accept, edit, or reject the result and continues the BI workflow. Meanwhile, a context management strategy [Uncaptioned image] automatically generates and maintains cell dependencies within the notebook to promote further agent calls. Next, we provide an overview of the three critical modules in DataLab.
@@
  • Inter-Agent Communication. This module formulates the information flow process among different agents as an FSM to enable more control over their communications, with nodes representing agents and edges representing inter-agent information transition directions. Upon task completion, each agent’s outputs are formatted into structured information units [18], comprising key characteristics such as the associated table’s identifier and a concise description of the agent’s executed actions. The module also maintains a shared information buffer for all agents to proactively exchange information based on the FSM to improve communication efficiency.
@@ IV-A Knowledge Generation
Traditionally, such knowledge has been manually constructed by domain experts, which is tedious and time-consuming. Through an extensive examination conducted at Tencent, it was observed that, while 85% of the tables lack comprehensive metadata, they are predominantly linked to various SQL or Python scripts utilized for data processing. These scripts reveal common usage patterns within practical business contexts. Additionally, for those tables lacking adequate processing scripts, data lineage information, which elucidates their connections to other tables or columns throughout the organization, provides an alternative resource for metadata imputation. Therefore, inspired by LLMs’ exceptional code understanding and reasoning abilities, we propose an LLM-based knowledge generation approach (Algorithm 1) that leverages script history to abstract and summarize knowledge components through meticulously-designed prompting techniques. This automated approach comprises a Map-Reduce process with a self-calibration mechanism [37] to generate high-quality knowledge for databases, tables, and columns.
@@
1:  // Filter out duplicated or similar scripts
2:  $\mathcal{H}\leftarrow\ preprocess(\mathcal{H})$
3:  Map Phase:
4:  $map\_res\leftarrow[\ ]$
5:  for each historical script $h_{i}\in\mathcal{H}$ do
6:     while $s_{i}<\mathcal{T}$ do
7:        // Generate knowledge individually
8:        $d_{i},t_{i},c_{i}\leftarrow\mathcal{LLM}(h_{i},\mathcal{S},\mathcal{L})$
9:        // Self-calibration
10:        $s_{i}\leftarrow\mathcal{LLM}(d_{i},t_{i},c_{i})$
11:     end while
12:     $map\_res.append([d_{i},t_{i},c_{i}])$
13:  end for
14:  Reduce Phase:
15:  // Synthesize knowledge collectively
16:  $\mathcal{D},\mathcal{T},\mathcal{C}\leftarrow\mathcal{LLM}(map\_res,\mathcal{S},\mathcal{L})$
17:  return  $\mathcal{D},\mathcal{T},\mathcal{C}$

### v3

# Yuhui Zhang, Yifeng Zheng, Xiuqi Huang, Minfeng Zhuparallel-$\parallel$, Yuxin Ma, Bin Cui§, Peng Chen‡ 🖂, Wei Chen† 🖂

@@ Abstract
Business intelligence (BI) transforms large volumes of data within modern organizations into actionable insights for informed decision-making. Recently, large language model (LLM)-based agents have streamlined the BI workflow by automatically performing task planning, reasoning, and actions in executable environments based on natural language (NL) queries. However, existing approaches primarily focus on individual BI tasks such as NL2SQL and NL2VIS. The fragmentation of tasks across different data roles and tools lead to inefficiencies and potential errors due to the iterative and collaborative nature of BI. In this paper, we introduce DataLab, a unified BI platform that integrates a one-stop LLM-based agent framework with an augmented computational notebook interface. DataLab supports various BI tasks for different data roles in data preparation, analysis, and visualization by seamlessly combining LLM assistance with user customization within a single environment. To achieve this unification, we design a domain knowledge incorporation module tailored for enterprise-specific BI tasks, an inter-agent communication mechanism to facilitate information sharing across the BI workflow, and a cell-based context management strategy to enhance context utilization efficiency in BI notebooks. Extensive experiments demonstrate that DataLab achieves state-of-the-art performance on various BI tasks across popular research benchmarks. Moreover, DataLab maintains high effectiveness and efficiency on real-world datasets from Tencent, achieving up to a 58.58% increase in accuracy and a 61.65% reduction in token cost on enterprise-specific BI tasks.
@@ I Introduction
Recent advancements in autonomous agents powered by large language models (LLMs) [3] offer the potential to streamline the BI workflow. By receiving instructions in natural language (NL), LLM-based agents can perform task planning, reasoning, and actions in executable environments. This can significantly reduce the complexity of many BI tasks, such as code generation [4], text-to-visualization translation [5], and automated insight discovery [6].
@@
C1: Lack of domain knowledge incorporation. Existing studies usually leverage clean and synthesized research benchmarks to build and evaluate agents [10]. However, BI tasks typically involve large and dirty real-world datasets with many ambiguities [11]. For example, column names in business data tables may have unclear semantic meanings [12], and user queries often include enterprise-specific jargon [10]. To mitigate these issues, incorporating extensive domain knowledge is essential to enhance agents’ understanding of input data and improve their performance on practical BI tasks. While some approaches adopt fine-tuning [13] or continued pre-training [14] to augment agents’ domain-specific capabilities, acquiring the necessary large and up-to-date training data corpora remains challenging in BI scenarios due to their complexity and dynamic nature. Other approaches (e.g., Chat2Data [15]) require users to manually integrate domain knowledge through external documents or customized knowledge bases, which is highly inefficient and inconvenient.
C2: Insufficient information sharing across tasks. Different tasks are typically managed by corresponding LLM-based agents to achieve optimal performance [16]. As a complex BI query may encompass multiple tasks, information sharing among the involved agents is critical. For example, the data retrieved by a SQL writing agent must be accurately conveyed to a chart generation agent. Therefore, an effective and efficient inter-agent communication mechanism is essential to align their understanding of the overall analysis goals, current data context, and executed actions. However, many existing multi-agent frameworks, such as ChatDev [17] and CAMEL [18], rely on unstructured natural language for communication, which can lead to distortions due to the inherent vagueness and redundancy of NL [19]. Consequently, they are inadequate for handling the complexity of BI tasks, which commonly involve diverse information types (e.g., data, charts, texts).
C3: Demand for adaptive LLM context management. LLM-based agents depend on their context windows (i.e., limited input tokens for NL understanding, reasoning, and generation) to complete tasks. Necessary contexts must be provided to ensure a successful and seamless workflow. Meanwhile, in a unified BI platform, vast amounts of multi-modal contexts (e.g., code snippets and their execution results, charts and their specifications) are intertwined and often relate to diverse data tables. Obviously, only relevant portions of these contexts are pertinent to specific tasks and should be selectively provided to the agents [20]. In contrast, existing works either focus on single-modal contexts tied to specific tasks [21] or indiscriminately provide all contexts [16], neither of which meets the demands of a unified BI platform. Thus, adaptive context management tailored for BI scenarios that considers prior states and current user needs is crucial for maintaining system efficiency and cost-effectiveness.
@@
To improve agents’ performance on enterprise-specific BI tasks (for C1), we develop a Domain Knowledge Incorporation module, a systematic approach for automated knowledge generation, organization, and utilization. It leverages data processing scripts (e.g., Python code, SQL queries) within the enterprise to extract knowledge associated with databases/tables/columns, thereby uncovering their common usage patterns. This module overcomes the practical challenge of manually integrating external knowledge bases for BI.
To facilitate information sharing across different tasks (for C2), we design an Inter-Agent Communication module, a structured mechanism that goes beyond pure NL to enhance the information representation capabilities of agents. It also formulates the information sharing process among agents with a finite state machine (FSM) for a more controlled and efficient flow of communication. This module addresses the challenge of effectively and efficiently sharing multi-modal information generated by various agents throughout the BI workflow.
To manage LLM contexts within multi-modal notebooks (for C3), we propose a Cell-based Context Management module that represents inter-cell dependencies using directed acyclic graphs (DAGs). These dependency graphs are dynamically updated in response to user modifications, enabling the adaptive selection of pertinent contexts based on specific task requirements. This module resolves the challenge of enhancing agents’ context utilization efficiency in BI scenarios.
Compared to existing works, DataLab stands out due to four benefits: (1) It delivers satisfactory performance across various BI tasks, rather than focusing on individual tasks; (2) It is well-suited for real-world applications, not just research benchmarks; (3) It offers a unified platform to satisfy different user requirements, rather than catering to a single data role; and (4) It integrates LLM assistance with user customization, instead of relying solely on end-to-end result generation.
@@
  • We extensively evaluate DataLab on both research benchmarks and real-world business datasets from Tencent, demonstrating its performance on various BI tasks. We also showcase the practical applications of DataLab at Tencent TEG’s Big Data Platform through user feedback.
@@ II-A BI Workflow
The BI workflow traditionally includes several key stages: data collection, storage, preparation, analysis, and visualization. Data Collection and Storage establish the foundation, typically supported by cloud infrastructures (e.g., Tencent TCHouse-C, AWS) and specialized data applications (e.g., BigQuery, Airflow) with mature ecosystem integration. Attempting to replicate these practices in alternative environments could introduce inefficiencies and anti-patterns. Moreover, the core value of modern BI lies in transforming raw data into actionable insights through data preparation, analysis, and visualization [1, 8]. Therefore, our work primarily focuses on the latter three stages, emphasizing interoperability with existing tools rather than reinventing them. For data collection and storage, we provide plugin APIs to enable seamless integration with third-party data connectors, ensuring flexibility and compatibility within the broader BI ecosystem.
Data Preparation [22] ensures data consistency, correctness, and quality. This usually includes cleaning, structuring, and enriching raw data into a format suitable for further analysis. Following preparation, Data Analysis [23] applies statistical and analytical techniques to extract insights, aiming to uncover patterns, trends, and correlations. Finally, Data Visualization [5] presents analyzed data in visual formats like charts, graphs, and dashboards, which makes complex data easier to understand and interpret for decision-makers.
The data roles involved in the BI workflow are specific to different organizations. Among them, data engineers, scientists, and analysts are usually indispensable. Data Engineers are primarily tasked with data preparation, constructing and administering data pipelines to ensure that data is accurately cleansed and structured for subsequent analysis. They typically use SQL and Python for data processing, and use cloud computing platforms like AWS for data storage and ingestion. Data Scientists engage in data preparation and analysis, applying advanced statistical and machine learning methodologies to extract insights and forecast trends from intricate datasets. They are familiar with Python/R and popular data science libraries like Pandas. Data Analysts concentrate on data analysis and visualization, analyzing data to discern patterns and conveying findings through detailed visual reports and dashboards. They use SQL to query data, and rely on BI platforms like Tableau to perform and share their analyses.
@@ II-B LLM-based Agents for BI
LLM-based agents are autonomous systems powered by LLMs that can perceive environments, execute tasks, make decisions, and interact with users in complex contexts [3]. These agents comprise profiling, memory, planning, and action modules, which respectively define the agent’s role, facilitate operations in dynamic environments through recall and future action planning, and convert decisions into outputs [24]. In BI scenarios, agents receive users’ NL queries and then perform data preparation, analysis, and visualization. By interpreting execution results, they can complete many BI tasks. For example, data preparation involves tasks like NL2SQL [25] and NL2DSCode [4], while data analysis and visualization rely on NL2Insight [26] and NL2VIS [5], respectively.
@@ III Overview
  • LLM-based Agent Framework. In DataLab, multiple agents are designed for different BI tasks based on user requirements. To achieve this, we first identify several common BI procedures and abstract them into data tools that can be called upon by agents during inference. Example tools include a Python sandbox for code execution and a Vega-Lite environment for visualization rendering. Accompanied by other auxiliary components like memory modules, each BI agent is represented as a DAG for high flexibility and easy extensibility. Within the DAG, nodes depict reusable components (e.g., LLM APIs, tools) and edges depict their connections (e.g., file transfer across tools). Figure 2 illustrates an example agent workflow for NL2VIS. Additionally, we add a proxy agent to the framework, which serves as a hub to directly interact with users and allocate tasks to each specialized agent based on user queries. Compared to existing approaches that primarily focus on individual tasks (e.g., NL2VIS [27], NL2SQL [21]), DataLab’s agent framework supports a wide array of tasks for various data roles across the BI workflow through multi-agent collaboration.
  • Computational Notebook Interface. DataLab’s notebook interface (Figure 3) serves as a unified, interactive, and collaborative environment for different data roles to complete their specialized tasks. To achieve this, we augment JupyterLab (a widely used notebook interface) to support (1) multi-language cells and (2) on-the-fly LLM assistance. First, DataLab wrangles SQL, Python/PySpark, Markdown, and Chart cells altogether, allowing both technical and non-technical users to easily adopt their familiar workflows on the notebook. Going beyond traditional notebooks that only support Python and Markdown, DataLab notebooks directly connect to backend databases for SQL query execution, and feature GUI-based dashboards [28] similar to Tableau for visualization authoring. Second, we integrate our LLM-based agent framework seamlessly into each notebook cell. Users can get LLM assistance both at notebook- and cell-level. Specifically, users toggle an input box and type their analytic queries, which are then processed by the agents in our framework. These agents can create new cells or modify existing ones in the notebook. Users can subsequently examine the results and make further customizations flexibly. Compared to existing end-to-end approaches (e.g., CHESS [29], LIDA [30]), DataLab’s notebook interface enables flexible user intervention to adapt LLM-generated results to real-world BI scenarios.
@@
DataLab Workflow. Upon receiving an NL query and the associated dataset, DataLab analyzes the dataset and interprets the query, incorporating domain knowledge (Figure 1[Uncaptioned image]) before feeding them into LLMs. Then, DataLab leverages various agents to complete the task, which may involve information sharing with each other through a structured communication mechanism (Figure 1[Uncaptioned image]). Subsequently, the corresponding result will be presented in the notebook. Users can either accept, edit, or reject the result and continues the BI workflow. Meanwhile, a context management strategy (Figure 1[Uncaptioned image]) automatically generates and maintains cell dependencies within the notebook to promote further agent calls. Next, we provide an overview of DataLab’s three critical modules.
@@
  • Inter-Agent Communication. This module formulates the information flow process among different agents as an FSM to enable more control over their communications, with nodes representing agents and edges representing inter-agent information transition directions. Upon task completion, each agent’s outputs are formatted into structured information units [19], comprising key characteristics such as the associated table’s identifier and a concise description of the actions performed. The module also maintains a shared information buffer to proactively exchange information based on the FSM to improve communication efficiency.
@@ IV-A Knowledge Generation
Many existing works assume that such knowledge can be manually constructed and integrated by domain experts. However, in large organizations, enormous numbers of data tables are commonly involved, making manual curation and maintenance highly impractical. To overcome this limitation, we first conducted an extensive investigation at Tencent. It was observed that, while 85% of the tables lack comprehensive metadata, they are predominantly linked to various SQL or Python scripts utilized for data processing. These scripts reveal common usage patterns within practical business contexts. Additionally, for those tables lacking adequate processing scripts, data lineage information, which elucidates their connections to other tables or columns throughout the organization, provides an alternative resource for metadata imputation. Therefore, inspired by LLMs’ exceptional code understanding and reasoning abilities, we propose an LLM-based knowledge generation approach (Algorithm 1) that leverages script history to abstract and summarize knowledge components through meticulously-designed prompting techniques. This automated approach comprises a Map-Reduce process with a self-calibration mechanism [32] to generate high-quality knowledge for databases, tables, and columns.
@@
1:  $\mathcal{H}\leftarrow\text{preprocess}(\mathcal{H})$ // Duplicated/Similar script filtering
[…]
3:  for each historical script $h_{i}\in\mathcal{H}$ do
4:     while $s_{i}<\mathcal{T}$ do
5:        $d_{i},t_{i},c_{i}\leftarrow\text{LLM}(h_{i},\mathcal{S},\mathcal{L})$ // Knowledge generation
6:        $s_{i}\leftarrow\text{LLM}(d_{i},t_{i},c_{i})$ // Self-calibration
7:     end while
[…]
9:  end for
10:  $\mathcal{D},\mathcal{T},\mathcal{C}\leftarrow\text{LLM}(\text{map\_res},\mathcal{S},\mathcal{L})$ // Knowledge synthesis
11:  return  $\mathcal{D},\mathcal{T},\mathcal{C}$